At a school play 129 tickets are sold. adult tickets cost a dollar and $.50 and the children cost 0.75 cents and In all $153 is taken and how many of each kind of ticket were sold?
Answer by precalc!!!!(2)

You must make two formula let x equal number of adult tickets and y equal number of children tickets each adult ticket is 1.5 so that mean its 1.5x and every children ticket was .75 so that means it was .75y
The total profit was $153 so then the formula must be 153= 1.5x + .75 y
Yet you cant work with two varibkes but you know that the number of adult and children at the show was 129 so: x+y=129 now solve for one varibke y= 129 -x and replace it in the equation before
153= 1.5x + .75(129-x)
153= 1.5x + 96.75 - .75x
153= .75x + 96.75
56,25=.75x
75= x
Now you know that 75 adult tickets were sold pulg that to the second equation
Y=129-x
Y=129-75
Y= 54
So 75 adult tickets were sold and 54 children tickets